OpenStudy (amtran_bus):

Is 1-x^2/2 a polynomial function.

OpenStudy (amtran_bus):

\[\frac{ 1-x^2 }{ 2 }\]

OpenStudy (anonymous):

A poynomial function is anything that has powers of x (and no other functions like sin,log,...) Roughly

OpenStudy (anonymous):

and exponents of x are integers

OpenStudy (amtran_bus):

But is the equation I gave one?

OpenStudy (amtran_bus):

Because It looks like a rational to me.

OpenStudy (anonymous):

What is the definition of a rational function?

OpenStudy (anonymous):

Its a rational if there is a function of x in the denominator (not here). If I do this to it\[-\frac{1}{2}x ^{2}+\frac{1}{2}\]then what do you think?

OpenStudy (amtran_bus):

Well, I guess it is a polynomial function due to the x^2 and no fractions.

OpenStudy (anonymous):

I think it is too.

OpenStudy (amtran_bus):

So a rational function has to have an x on the denominator. And a poly. does not.

OpenStudy (anonymous):

I would say yes.

OpenStudy (amtran_bus):

I like school buses. Thanks!

OpenStudy (anonymous):

Me too! woof!